# Improving the Cosmological Constraints by Inferring the Formation Channel of Extreme-mass-ratio Inspirals

### v1

@@ Abstract
Extreme-mass ratio inspirals (EMRIs) could be detected by space-borne gravitational-wave (GW) detectors, such as the Laser Interferometer Space Antenna (LISA), TianQin and Taiji. In general, locating EMRIs by space-borne GW detectors can help us select the candidate host galaxies which can be used to infer the cosmic expansion history. In this paper, we show that the localization information of EMRIs can also be used to select the candidate host AGNs which can be used to infer the formation channel of the EMRIs and extract more precisely the redshift probability distributions of the EMRIs. Using the EMRIs that are expected to be detected by TianQin and LISA and the galaxy catalog that can be provided by the Chinese Space Station Telescope for the analysis, we find that TianQin can constrain the Hubble-Lemaître constant $H_{0}$ to a precision of about $3\%-8\%$ and the dark energy equation of state parameter $w_{0}$ to about $10\%-40\%$, the network composed of TianQin and LISA can improve the precisions of $H_{0}$ and $w_{0}$ to about $0.4\%-7\%$ and $4\%-20\%$, respectively, without considering the effects of AGNs. Furthermore, combining the detected EMRIs with AGN catalog by the statistical framework of likelihood-ratio-based, we find TianQin can establish the EMRI-AGN correlation with about $500$ EMRIs if all of the EMRIs indeed originate in AGNs, and the TianQin+LISA network can reduce this required number to about $30$. Once the EMRI-AGN correlation is confirmed, the combination of the EMRIs and AGN catalog can significantly improve the constraints on the cosmological parameters. These results demonstrate the potentials of using EMRIs as well as galaxy and AGN catalogs to constrain the cosmological parameters and astrophysical models.
@@
software: numpy (van der Walt et al., 2011), scipy (Virtanen et al., 2020), emcee (Foreman-Mackey et al., 2013, 2019), matplotlib (Hunter, 2007), corner (Foreman-Mackey, 2016), GetDist (Lewis, 2019), and seaborn (Waskom, 2021). thanks: Boya fellow
@@ 1 Introduction
The first three observing runs of the Advanced LIGO and Virgo, and KAGRA detectors (LVK) have detected about $90$ gravitational-wave (GW) events triggered by the mergers of compact-object binaries (Abbott et al., 2016, 2017a, 2019, 2021b, 2023a). In addition, increasing evidence for the existence of GW signal in the nanohertz band is accumulating thanks to the observations of pulsar timing arrays (Agazie et al., 2023; Antoniadis et al., 2023; Reardon et al., 2023; Xu et al., 2023), although the origin of this GW signal has yet to be identified. These observations opened a new era of observational astronomy, in which GWs can be used as new probes of fundamental physics (Abbott et al., 2021c, d), astrophysics (Fishbach & Holz, 2017; Abbott et al., 2020b; Fishbach et al., 2021; Abbott et al., 2023b; Bartos et al., 2017; Veronesi et al., 2022, 2023), and the cosmic expansion history (Abbott et al., 2017b; Soares-Santos et al., 2019; Fishbach et al., 2019; Hotokezaka et al., 2019; Palmese et al., 2020; Vasylyev & Filippenko, 2020; Abbott et al., 2021a; Wang & Giannios, 2021; Finke et al., 2021; Abbott et al., 2023c).
The idea of combining a GW event and its electromagnetic (EM) counterpart to establish a “standard siren” which can be used to probe the expansion history of the Universe, such as measuring the Hubble-Lemaître constant, was first proposed by Schutz (1986). Many researchers followed this idea and have established the methodological framework even before the detection of GW (Marković, 1993; Holz & Hughes, 2005; Dalal et al., 2006; Arun et al., 2009; Petiteau et al., 2011; Babak et al., 2011; Taylor et al., 2012; Del Pozzo, 2012; Nissanke et al., 2013). The first successful measurement of the Hubble-Lemaître constant using standard siren was accomplished in 2017 (Abbott et al., 2017b) immediately after the detection of a binary neutron star merger in both GWs (Abbott et al., 2017a) and EM waves (Abbott et al., 2017c). Even when a unique EM counterpart is difficult to identify, the redshift may be inferred from the statistical distribution of the galaxies in the same sky area localized by GW detectors (Schutz, 1986; MacLeod & Hogan, 2008; Petiteau et al., 2011; Del Pozzo, 2012). Such “dark sirens” also provide important constraints on the Hubble-Lemaître constant, e.g., in the cases of GW170814 (Soares-Santos et al., 2019) and GW190814 (Palmese et al., 2020; Vasylyev & Filippenko, 2020).
@@
The third type of low-frequency GW sources which can be used as standard sirens are EMRIs. EMRI normally resides at the center of a galaxy and consists of an MBH and an inspiralling stellar-mass BH (Amaro-Seoane, 2018). The potential of using LISA EMRIs to study cosmology was first pointed out in MacLeod & Hogan (2008), and later studied in more detail by the literatures (Laghi et al., 2021; Liu et al., 2023). These earlier studies considered only the EMRIs with $z<1$ where the galaxy catalogs are relatively more complete. However, population studies have shown that space-borne GW detectors can catch EMRIs up to $z\approx 3$ or higher (Babak et al., 2017; Fan et al., 2020; Bonetti & Sesana, 2020). At such high redshifts, although the current galaxy catalog is incomplete, future space telescopes will be able to conduct a more complete survey of the galaxies (Gong et al., 2019; Ivezić et al., 2019; Scaramella et al., 2022). In addition, beside LISA, other space-borne GW detectors, such as TianQin and Taiji, can also detect EMRIs. A joint observation by multiple detectors in principle can improve the sky localization of a GW source by orders of magnitude (Ruan et al., 2020; Gong et al., 2021). These observational prospects in general should improve the constraint on cosmological parameters by EMRIs.
@@
The prospect of using EMRIs to constrain the cosmological parameters relies on not only the completeness of a galaxy catalog, but also the accuracy of the redshift which can be inferred from the galaxy catalog. When extracting the redshift probability distributions of EMRIs from galaxy catalogs, previous works (MacLeod & Hogan, 2008; Laghi et al., 2021; Liu et al., 2023) have assumed that all galaxies have an equal probability to host an EMRI. However, in theory EMRIs could form in either normal galaxies (Amaro-Seoane, 2018) or active galactic nuclei (AGNs, e.g. Yunes et al., 2011), and recent models suggest that the latter channel could be more important because the accretion disks in AGNs can capture and drive a large number of stellar-mass BHs to migrate towards the center (Pan et al., 2021). Therefore, what type of galaxies in a catalog can be considered as the hosts of EMRIs remains a question.
@@
Recently, a statistical framework has been proposed which allows one to test the spatial correlation between GW sources and AGNs even in the absence of identifying the host galaxies of the GW sources Bartos et al. (2017). The framework utilizes a Poisson distribution to describe the likelihood of the number of candidate host AGNs for the GW source, and constrains the spatial correlation between the GW sources and AGNs by a likelihood-ratio-based method. Several works have demonstrated the effectiveness of this framework in identifying the host galaxies of the LVK binary BHs (Veronesi et al., 2022, 2023) or the MBHBs of future space-borne GW detectors (Zhu & Chen, 2024). Such a method in principle can be used to infer the fraction of EMRIs which originate in AGNs. Knowing this fraction will help us better weigh the large number of candidate host galaxies in the sky area localized for an EMRI, so that we can more precisely extract the redshift probability distributions and measure cosmological parameters.
@@
Therefore, in this work we will apply the aforementioned statistical framework to EMRIs, and study its effectiveness in constraining the cosmological parameters and testing the spatial correlation between EMRIs and AGNs. To facilitate the study, we consider TianQin, as well as a network composed of TianQin and LISA, as the future observational instruments. This paper is organized as follows. In Section 2 we introduce the statistical analysis framework that we use to infer the cosmic expansion history and test the EMRI-AGN correlation. In Section 3 we conduct mock observations. The results are shown in Section 4 for the cosmological prospects and Section 5 for the prospects of inferring the EMRI-AGN correlation. In Section 6, we discuss some of the factors that can affect the resuls. Finally, in Section 7 we summarize the main findings of this work and the implications of our results.
@@ 2.1 Cosmological model
In this work, we consider two dark energy models, namely the standard model—the cosmological constant $\Lambda$, and the Chevallier-Polarski-Linder (CPL) model (Chevallier & Polarski, 2001; Linder, 2003). (i) For the cosmological constant model, the EoS of the dark energy is constantly equal to minus one, i.e., $w(z)\equiv-1$. This model plus the cold dark matter model, plus the spatially-flat assumption, make up the so-called standard cosmological model—flat-$\Lambda$CDM. (ii) For the CPL model, the EoS of the dark energy is modeled as a function, $w(z)=w_{0}+w_{a}{z}/{(1+z)}$ (Chevallier & Polarski, 2001), here $w_{0}$ and $w_{a}$ are the extra parameters. The analytic expression of Equation (1) is $H(z)=H_{0}\sqrt{\Omega_{M}(1\!+\!z)^{3}+\Omega_{\Lambda}(1\!+\!z)^{3(1+w_{0}+w_{a})}\exp\!\big{(}\!\!-\!\frac{3w_{a}z}{1+z}\!\big{)}}$. In particular, when $w_{0}=-1$ and $w_{a}=0$, the CPL model will degenerate with the cosmological constant model.
@@
where $c$ is the speed of light. In our simulations, we fix all the cosmological parameters to the values: $H_{0}\equiv h\times 100~{}\!{\rm km}~{}\!{\rm s}^{-1}~{}\!\!{\rm Mpc}^{-1}=67.8~{}\!{\rm km}~{}\!{\rm s}^{-1}~{}\!\!{\rm Mpc}^{-1}$ (here $h$ is the dimensionless Hubble-Lemaître constant), $\Omega_{M}=0.307$, $w_{0}=-1$, and $w_{a}=0$.
@@ 2.2 Bayesian framework for GW cosmology
To constrain the cosmological parameters by combining the GW and EM data, we adopt the Bayesian framework presented in Chen et al. (2018). However the estimations of the completeness of the galaxy catalogs were modified according to the correlation between the EMRIs’ MBHs and their host galaxies, such as the approximate linear relationship between the masses of MBHs and the (bulge) luminosities of the galaxies (which is known as the $M_{\rm MBH}-L$ relation; Ferrarese & Merritt, 2000; Scott & Graham, 2013; Kormendy & Ho, 2013).
@@
$\frac{\int\!p(d^{\rm gw}_{i}\!,\!d^{\rm em}_{i}\!,\!\alpha,\!\delta,\!D_{L},\!z,\!M,\!L|{\bf\Omega},\!I)\mathrm{d}\alpha\mathrm{d}\delta\mathrm{d}D_{L}\mathrm{d}z\mathrm{d}M\mathrm{d}L}{\beta({\bf\Omega}|I)},$ (4)
where $\beta({\bf\Omega}|I)$ is a normalization factor that use for accounts for the selection effects in GW and EM observations (Mandel et al., 2019; Chen et al., 2018; Soares-Santos et al., 2019).
@@
$~{}p(d^{\rm gw}_{i}|\alpha,\delta,D_{L},M,{\bf\Omega},I)p(d^{\rm em}_{i}|\alpha,\delta,z,L,{\bf\Omega},I)$
@@
The term $p(d^{\rm gw}_{i}|\alpha,\delta,D_{L},M,{\bf\Omega},I)$ is the likelihood of the GW data (Finn, 1992; Sathyaprakash & Schutz, 2009) after marginalized over all other parameters that unrelated to the cosmic expansion history, and $p(d^{\rm em}_{i}|\alpha,\delta,z,L,{\bf\Omega},I)$ is the likelihood of the corresponding EM data. For dark sirens, because there is no information from the observations of the EM counterparts, we can define $d^{\rm em}_{i}$ to null, as well as define the likelihood $p(d^{\rm em}_{i}|\alpha,\delta,z,L,{\bf\Omega},I)$ to constant (Chen et al., 2018). The term $p_{0}(D_{L}|z,{\bf\Omega},I)$ is a Dirac delta function, i.e., $p_{0}(D_{L}|z,{\bf\Omega},I)=\delta_{\rm D}\big{(}D_{L}-\hat{D}_{L}(z,{\bf\Omega})\big{)}$ (here $\hat{D}_{L}(z,{\bf\Omega})$ represents Equation (2)), because $D_{L}$ is a function of $z$ given a cosmology.
@@
where the two values in square bracket $\mathcal{U}[...~{}\!,...]$ represent the lower and upper boundaries of the uniform distribution, $\hat{M}(L)$ represents the mass of the MBH as a function of the luminosity of the host galaxy, and $\sigma_{\!M}$ is the standard deviation. The expression of $p_{0}(M|z,L,{\bf\Omega},I)$ holds just for EMRI (and MBHB) GW sources due to the $M_{\rm MBH}-L$ relation, the reason for adopting a uniform rather than a Gaussian distribution is that the data for which the $M_{\rm MBH}-L$ relation is derived have many outliers (Ferrarese & Merritt, 2000; Scott & Graham, 2013; Kormendy & Ho, 2013). The function $\hat{M}(L)$ can be written as
@@
(Scott & Graham, 2013), where $a_{\rm fit}$ and $b_{\rm fit}$ are the fitting zeropoint and slope, and $M_{\odot}\!=\!1.989\times 10^{30}~{}\!{\rm kg}$, $M^{*}_{\odot}\!=\!+4.8\!~{}{\rm mag}$ and $L_{\odot}\!=\!3.8\times 10^{33}\!~{}{\rm erg~{}s}^{-1}$ are the mass, absolute magnitude, and luminosity of the Sun, respectively. The best-fitting parameters for Equation (7) in the $K$-band are $a_{\rm fit}=8.04$ and $b_{\rm fit}=-0.48$ with a root-mean-square scatter of $\sigma_{M}=0.4$ in $\lg(M/M_{\odot})$ direction (Scott & Graham, 2013). The reason, moreover, that the uniform distribution we adopt is not symmetric around $\hat{M}(L)$ is that, there is a fraction of galaxies with lighter MBHs compared to the predicted $\hat{M}(L)$ (Jiang et al., 2011; Scott & Graham, 2013; Kormendy & Ho, 2013).
The last term $p_{0}(\alpha,\delta,z,L|{\bf\Omega},I)$ in Equation (2.2) is a prior distribution on the spatial positions and luminosities of galaxies. To get a proper prior, we adopt the standard assumption that the spatial probability distribution of compact binary GW sources is proportional to the number density of galaxies (Schutz, 1986; MacLeod & Hogan, 2008; Petiteau et al., 2011; Del Pozzo, 2012; Chen et al., 2018; Fishbach et al., 2019; Soares-Santos et al., 2019; Gray et al., 2020; Finke et al., 2021; Abbott et al., 2023c), while taking into account the selection effects that many galaxies are too dim to be recorded by surveys, we can express the prior as
@@
where $p_{\rm obs}(\alpha,\delta,z,L|{\bf\Omega},I)$ represents a distribution function of observed galaxies from EM surveys, $f_{\rm compl}$ is the completeness fraction of the observed galaxy catalog, and $p_{\rm miss}(\alpha,\delta,z,L|{\bf\Omega},I)$ represents the probability distribution function of the unobservable galaxies.
@@
where $N_{\rm gal}$ is the total number of the observed galaxies and two $\mathcal{N}[\bar{x},\sigma_{x}](x)$ represent Gaussian distributions of $x$ with mean values $\bar{x}$ and standard deviations $\sigma_{x}$. Here, because the errors of sky positions of galaxies in EM surveys are much smaller than the sky localization errors of the GW sources, the distribution of $(\alpha,\delta)$ of galaxies can be approximately expressed as two Dirac delta functions. (ii) For the unobservable galaxies, their probability distribution function can be expressed as
@@
is the complementary error function with mean value $\hat{L}_{\rm thr}(z,{\bf\Omega})$ and standard distribution $\bar{\sigma}_{L}$, $\Phi(L,z)$ represents the luminosity function of the galaxies at $z$ (Schechter, 1976; Dahlen et al., 2005; Marchesini et al., 2007), $V_{\rm c}$ is the comoving volume, and $\mathrm{d}\hat{\Omega}=\cos\delta\!~{}\mathrm{d}\alpha\mathrm{d}\delta$ represents the differential solid angle. The mean value $\hat{L}_{\rm thr}(z,{\bf\Omega})$ is a luminosity threshold for an observable galaxy with $z$ in a cosmology ${\bf\Omega}$, it depends on the limit magnitude $m_{\rm limit}^{*}$ of the galaxy survey, we can calculate it with
$\!\!\!\!\!\!\!\!\!\!\lg\!\left(\!\frac{\hat{L}_{\rm thr}(z,{\bf\Omega})}{L_{\odot}}\!\right)\!=$ $*\displaystyle~{}\!\frac{1}{2.5}\Bigg{[}M^{*}_{\odot}-m_{\rm limit}^{*}$
@@
Finally, to calculate the denominator $\beta({\bf\Omega}|I)$ in Equation (2.2), we adopt the same estimated approach as presented in Chen et al. (2018) (also see Gray et al., 2020; Finke et al., 2021; Abbott et al., 2023c; Zhu & Chen, 2023). For dark sirens, the denominator is determined by the GW data $d^{\rm gw}$ and the prior $p_{0}(\alpha,\delta,z,L|{\bf\Omega},I)$, we have
@@
Here, the equation represents to integrate over the possible GW data $d^{\rm gw}$ that the signal-to-noise ratio (SNR) is greater than the detection threshold.

### v2

@@ Abstract
Extreme-mass-ratio inspirals (EMRIs) could be detected by space-borne gravitational-wave (GW) detectors, such as the Laser Interferometer Space Antenna (LISA), TianQin and Taiji. Localizing EMRIs by GW detectors can help us select candidate host galaxies, which can be used to infer the cosmic expansion history. In this paper, we demonstrate that the localization information can also be used to infer the formation channel of EMRIs, and hence allow us to extract more precisely the redshift probability distributions. By conducting mock observations of the EMRIs which can be detected by TianQin and LISA, as well as the galaxies which can be provided by the future Chinese Space Station Telescope, we find that TianQin can constrain the Hubble-Lemaître constant $H_{0}$ to a precision of $\sim 3\%-8\%$ and the dark energy equation of state parameter $w_{0}$ to $\sim 10\%-40\%$. The TianQin+LISA network, by increasing the localization accuracy, can improve the precisions of $H_{0}$ and $w_{0}$ to $\sim 0.4\%-7\%$ and $\sim 4\%-20\%$, respectively. Then, considering an illustrative case in which all EMRIs originate in AGNs, and combining the mock EMRI observation with a mock AGN catalog, we show that TianQin can recognize the EMRI-AGN correlation with $\sim 1300$ detections. The TianQin+LISA network can reduce this required number to $\sim 30$. Additionally, we propose a statistical method to directly estimate the fraction of EMRIs produced in AGNs, $f_{\rm agn}$, and show that observationally deriving this value could significantly improve the constraints on the cosmological parameters. These results demonstrate the potentials of using EMRIs as well as galaxy and AGN surveys to improve the constraints on cosmological parameters and the formation channel of EMRIs.
@@
software: numpy (van der Walt et al., 2011), scipy (Virtanen et al., 2020), emcee (Foreman-Mackey et al., 2013, 2019), matplotlib (Hunter, 2007), corner (Foreman-Mackey, 2016), GetDist (Lewis, 2019), and seaborn (Waskom, 2021).
{CJK*}
UTF8gbsn
thanks: Boya fellow
@@ 1 Introduction
The first three observing runs of the Advanced LIGO and Virgo, and KAGRA detectors (LVK) have detected about $90$ gravitational-wave (GW) events triggered by the mergers of compact-object binaries (Abbott et al., 2016, 2017a, 2019a, 2021b, 2023a). In addition, evidence for the existence of GW signal in the nanohertz band is accumulating thanks to the observations of pulsar timing arrays (Agazie et al., 2023; Antoniadis et al., 2023; Reardon et al., 2023; Xu et al., 2023). These observations opened a new era of observational astronomy, in which GW can be used as a new probe of fundamental physics (Abbott et al., 2021c, d), astrophysics (Fishbach & Holz, 2017; Abbott et al., 2020b; Fishbach et al., 2021; Abbott et al., 2023b; Bartos et al., 2017; Veronesi et al., 2022, 2023), and the cosmic expansion history (Abbott et al., 2017b; Soares-Santos et al., 2019; Fishbach et al., 2019; Hotokezaka et al., 2019; Palmese et al., 2020; Vasylyev & Filippenko, 2020; Abbott et al., 2021a; Wang & Giannios, 2021; Finke et al., 2021; Abbott et al., 2023c).
The idea of combining a GW event and its electromagnetic (EM) counterpart to establish a “standard siren” to probe the expansion history of the universe, such as measuring the Hubble-Lemaître constant, was first proposed by Schutz (1986). The idea was followed by many later works which gradually established a methodological framework even before the detection of GW (Marković, 1993; Holz & Hughes, 2005; Dalal et al., 2006; Arun et al., 2009; Petiteau et al., 2011; Babak et al., 2011; Taylor et al., 2012; Del Pozzo, 2012; Nissanke et al., 2013). The first successful measurement of the Hubble-Lemaître constant using standard siren was accomplished in 2017 (Abbott et al., 2017b) immediately after the detection of a binary neutron star merger in both GWs (Abbott et al., 2017a) and EM waves (Abbott et al., 2017c). For other GW events, however, a unique EM counterpart is difficult to identify, and hence the redshift should be inferred from the statistical distribution of the galaxies in the same sky area localized by GW detectors (Schutz, 1986; MacLeod & Hogan, 2008; Petiteau et al., 2011; Del Pozzo, 2012). Such “dark sirens” also provide important constraints on the Hubble-Lemaître constant, e.g., in the cases of GW170814 (Soares-Santos et al., 2019) and GW190814 (Palmese et al., 2020; Vasylyev & Filippenko, 2020).
@@
The third type of low-frequency GW sources which can be used as standard sirens are EMRIs. EMRI normally resides at the center of a galaxy and consists of an MBH and an inspiralling stellar-mass BH (Amaro-Seoane, 2018). The potential of using LISA EMRIs to study cosmology was first pointed out in MacLeod & Hogan (2008), and later studied in more detail in Laghi et al. (2021); Liu et al. (2023). These earlier studies considered only the EMRIs with $z<1$ where the galaxy catalogs are relatively more complete. However, population studies have shown that space-borne GW detectors can catch EMRIs up to $z\approx 3$ or higher (Babak et al., 2017; Fan et al., 2020; Bonetti & Sesana, 2020). At such high redshifts, although the current galaxy catalog is incomplete, future space telescopes will be able to conduct a more complete survey of the galaxies (Gong et al., 2019; Ivezić et al., 2019; Scaramella et al., 2022). In addition, besides LISA, other space-borne GW detectors, such as TianQin and Taiji, can also detect EMRIs. A joint observation by multiple detectors in principle can improve the sky localization of a GW source by orders of magnitude (Ruan et al., 2020; Gong et al., 2021). These observational prospects in general should improve the constraint on cosmological parameters by EMRIs.
@@
The prospect of using EMRIs to constrain the cosmological parameters relies on not only the completeness of a galaxy catalog, but also the accuracy of the redshift which can be inferred from the galaxy catalog. When extracting the redshift probability distributions of EMRIs from galaxy catalogs, previous works (MacLeod & Hogan, 2008; Laghi et al., 2021; Liu et al., 2023) have assumed that all galaxies have an equal probability to host an EMRI. However, there might exist a substantial bias towards specific types of galaxies in the formation of EMRIs, as alternative formation channels could potentially yield higher EMRI rates in comparison to the standard channel (Hopman & Alexander, 2005; Babak et al., 2017; Amaro-Seoane, 2018). These alternative channels include binary separations (Miller et al., 2005), MBHB systems (Mazzolari et al., 2022; Naoz et al., 2022; Naoz & Haiman, 2023) and AGNs (Levin, 2003, 2007; Yunes et al., 2011; Pan & Yang, 2021; Pan et al., 2021; Derdzinski & Mayer, 2023). In particular, AGNs are potentially favourable environments for forming EMRIs because the accretion disks of AGNs can capture more stellar-mass BHs and BH binaries and drive them to migrate towards the central MBHs than normal galactic nuclei can do (Tagawa et al., 2020; Abbott et al., 2023b). The major properties predicted by the different EMRI formation channels, such as the EMRI rates, eccentricities at plunges, and masses of the MBHs and stellar-mass BHs, are subject to considerable uncertainties, posing challenges in determining the origins of EMRIs. Therefore, what type of galaxies in a catalog can be considered as the hosts of EMRIs remains a question.
@@
Recently, a statistical framework has been proposed which allows one to test the spatial correlation between GW sources and AGNs even in the absence of identifying the host galaxies of the GW sources (Bartos et al., 2017). The framework utilizes a Poisson distribution to describe the likelihood of the number of candidate host AGNs for the GW source, and constrains the spatial correlation between the GW sources and AGNs by a likelihood-ratio-based method. Several works have demonstrated the effectiveness of this framework in identifying the host galaxies of the LVK binary BHs (Veronesi et al., 2022, 2023) or the MBHBs of future space-borne GW detectors (Zhu & Chen, 2024). Such a method in principle can be used to infer the fraction of EMRIs which originate in AGNs. Knowing this fraction will help us better weigh the large number of candidate host galaxies in the sky area localized for an EMRI, so that we can more precisely extract the redshift probability distributions and measure cosmological parameters.
@@
Therefore, in this work we will apply the aforementioned statistical framework to EMRIs, and study its effectiveness in constraining the cosmological parameters and testing the spatial correlation between EMRIs and AGNs. To facilitate the study, we consider TianQin, as well as a network composed of TianQin and LISA, as the future observational instruments. This paper is organized as follows. In Section 2 we introduce the statistical analysis framework that we use to infer the cosmic expansion history and test the EMRI-AGN correlation. In Section 3 we conduct mock observations. The results are shown in Section 4 for the cosmological prospects and Section 5 for the prospects of inferring the EMRI-AGN correlation. In Section 6, we discuss some of the factors that can affect the results. Finally, in Section 7 we summarize the main findings of this work and the implications of our results.
@@ 2.1 Cosmological model
In this work, we consider two dark energy models, namely the standard model—the cosmological constant $\Lambda$, and the Chevallier-Polarski-Linder (CPL) model (Chevallier & Polarski, 2001; Linder, 2003). (i) For the cosmological constant model, the EoS of the dark energy is constantly equal to minus one, i.e., $w(z)\equiv-1$. This model, together with the cold dark matter model and the spatially-flat assumption, constitute the so-called standard cosmological model—flat-$\Lambda$CDM. (ii) For the CPL model, the EoS of the dark energy is modeled as a function of $w(z)=w_{0}+w_{a}{z}/{(1+z)}$ (Chevallier & Polarski, 2001), where $w_{0}$ and $w_{a}$ are two extra parameters. The analytic expression of Equation (1) is $H(z)=H_{0}\sqrt{\Omega_{M}(1\!+\!z)^{3}+\Omega_{\Lambda}(1\!+\!z)^{3(1+w_{0}+w_{a})}\exp\!\big{(}\!\!-\!\frac{3w_{a}z}{1+z}\!\big{)}}$. In particular, when $w_{0}=-1$ and $w_{a}=0$, the CPL model recovers the cosmological constant model.
@@
where $c$ is the speed of light. In our simulations, we fix the cosmological parameters to the values: $H_{0}\equiv h\times 100~{}\!{\rm km}~{}\!{\rm s}^{-1}~{}\!\!{\rm Mpc}^{-1}=67.8~{}\!{\rm km}~{}\!{\rm s}^{-1}~{}\!\!{\rm Mpc}^{-1}$ (here $h$ is the dimensionless Hubble-Lemaître constant), $\Omega_{M}=0.307$, $w_{0}=-1$, and $w_{a}=0$.
@@ 2.2 Bayesian framework for GW cosmology
To constrain the cosmological parameters by combining the GW and EM data, we adopt the Bayesian framework presented in Chen et al. (2018). However the estimation of the completeness of the galaxy catalog is modified. Here the estimation is based on the correlation between the MBHs in EMRIs and their host galaxies. More specifically, we use the approximate linear relationship between the masses of MBHs and the (bulge) luminosities of the galaxies (which is known as the $M_{\rm MBH}-L$ relation; Ferrarese & Merritt, 2000; Scott & Graham, 2013; Kormendy & Ho, 2013).
@@
[…]
where $\beta({\bf\Omega}|I)$ is a normalization factor accounting for the selection effects in GW and EM observations (Mandel et al., 2019; Chen et al., 2018; Soares-Santos et al., 2019).
@@
[…]
@@
The term $p(d^{\rm gw}_{i}|\alpha,\delta,D_{L},M,{\bf\Omega},I)$ is the likelihood of the GW data (Finn, 1992; Sathyaprakash & Schutz, 2009) after marginalized over all other parameters that are unrelated to the cosmic expansion history, and $p(d^{\rm em}_{i}|\alpha,\delta,z,L,{\bf\Omega},I)$ is the likelihood of the corresponding EM data. For dark sirens, because there is no information from the observations of the EM counterparts, we can define $d^{\rm em}_{i}$ to null, as well as define the likelihood $p(d^{\rm em}_{i}|\alpha,\delta,z,L,{\bf\Omega},I)$ to constant (Chen et al., 2018). The term $p_{0}(D_{L}|z,{\bf\Omega},I)$ is a Dirac delta function, i.e., $p_{0}(D_{L}|z,{\bf\Omega},I)=\delta_{\rm D}\big{(}D_{L}-\hat{D}_{L}(z,{\bf\Omega})\big{)}$ (here $\hat{D}_{L}(z,{\bf\Omega})$ represents Equation (2)), because $D_{L}$ is a function of $z$ given a cosmology.
@@
where the two values in square bracket $\mathcal{U}[...~{}\!,...]$ represent the lower and upper boundaries of the uniform distribution, $\hat{M}(L)$ represents the mass of the MBH as a function of the luminosity of the host galaxy, and $\sigma_{\!M}$ is the standard deviation. The expression of $p_{0}(M|z,L,{\bf\Omega},I)$ holds just for EMRI (and MBHB) GW sources due to the $M_{\rm MBH}-L$ relation. The reason for adopting a uniform rather than a Gaussian distribution is that the data for which the $M_{\rm MBH}-L$ relation is derived have many outliers (Ferrarese & Merritt, 2000; Scott & Graham, 2013; Kormendy & Ho, 2013). The function $\hat{M}(L)$ can be written as
@@
(Scott & Graham, 2013), where $a_{\rm fit}$ and $b_{\rm fit}$ are the fitting zeropoint and slope, and $M_{\odot}\!=\!1.989\times 10^{30}~{}\!{\rm kg}$, $M^{*}_{\odot}\!=\!+4.8\!~{}{\rm mag}$ and $L_{\odot}\!=\!3.8\times 10^{33}\!~{}{\rm erg~{}s}^{-1}$ are the mass, absolute magnitude, and luminosity of the Sun, respectively. The best-fitting parameters for Equation (7) in the $K$-band are $a_{\rm fit}=8.04$ and $b_{\rm fit}=-0.48$ with a root-mean-square scatter of $\sigma_{M}=0.4$ in $\lg(M/M_{\odot})$ (Scott & Graham, 2013). Notice that the uniform distribution adopted by us is not symmetric around $\hat{M}(L)$ because there is a fraction of galaxies with lighter MBHs compared to the predicted $\hat{M}(L)$ (Jiang et al., 2011; Scott & Graham, 2013; Kormendy & Ho, 2013).
The last term $p_{0}(\alpha,\delta,z,L|{\bf\Omega},I)$ in Equation (2.2) is a prior distribution for the spatial positions and luminosities of galaxies. To get a proper prior, we adopt the standard assumption that the spatial probability distribution of GW sources is proportional to the number density of galaxies (Schutz, 1986; MacLeod & Hogan, 2008; Petiteau et al., 2011; Del Pozzo, 2012; Chen et al., 2018; Fishbach et al., 2019; Soares-Santos et al., 2019; Gray et al., 2020; Finke et al., 2021; Abbott et al., 2023c). Further taking into account the selection effect that many galaxies are too dim to be recorded by surveys, we can express the prior as
@@
where $p_{\rm obs}(\alpha,\delta,z,L|{\bf\Omega},I)$ represents a distribution function of observed galaxies from EM surveys, $f_{\rm compl}$ is the completeness fraction of the observed galaxy catalog, and $p_{\rm miss}(\alpha,\delta,z,L|{\bf\Omega},I)$ represents the probability distribution function (PDF) of the unobservable galaxies.
@@
where $N_{\rm gal}$ is the total number of the observed galaxies and two $\mathcal{N}[\bar{x},\sigma_{x}](x)$ represent Gaussian distributions of $x$ with mean values $\bar{x}$ and standard deviations $\sigma_{x}$. Here, because the errors of sky positions of galaxies in EM surveys are much smaller than the sky localization errors of the GW sources, the distribution of $(\alpha,\delta)$ of galaxies can be approximately expressed as two Dirac delta functions. (ii) For the unobservable galaxies, their PDF can be expressed as
@@
is the complementary error function with mean value $\hat{L}_{\rm thr}(z,{\bf\Omega})$ and standard distribution $\bar{\sigma}_{L}$, $\Phi(L,z)$ represents the luminosity function of the galaxies at $z$ (Schechter, 1976; Dahlen et al., 2005; Marchesini et al., 2007), $V_{\rm c}$ is the comoving volume, and $\mathrm{d}\hat{\Omega}=\cos\delta\!~{}\mathrm{d}\alpha\mathrm{d}\delta$ represents the differential solid angle. The mean value $\hat{L}_{\rm thr}(z,{\bf\Omega})$ is a luminosity threshold for an observable galaxy with $z$ in a cosmology ${\bf\Omega}$, and it depends on the limiting magnitude $m_{\rm limit}^{*}$ of the galaxy survey. We can calculate it with
$\!\!\!\!\!\!\!\!\!\!\lg\!\left(\!\frac{\hat{L}_{\rm thr}(z,{\bf\Omega})}{L_{\odot}}\!\right)\!=$ $~{}\!\frac{1}{2.5}\Bigg{[}M^{*}_{\odot}-m_{\rm limit}^{*}$
@@
Finally, to calculate the denominator $\beta({\bf\Omega}|I)$ in Equation (2.2), we adopt the same estimated approach as presented in Chen et al. (2018) (also see Gray et al., 2020; Finke et al., 2021; Abbott et al., 2023c; Zhu & Chen, 2023). For dark sirens, the denominator is determined by the GW data $d^{\rm gw}$ and the prior $p_{0}(\alpha,\delta,z,L|{\bf\Omega},I)$, and we have
@@
Here, the equation represents an integration over the possible GW data $d^{\rm gw}$ that the signal-to-noise ratio (SNR) is greater than the detection threshold.